1. the grams of water that must be added to 4.00 g KI in the preparation of 2.00 % KI solution.


1
Expert's answer
2021-10-27T06:45:25-0400

Mass of solution = mass of solvent+mass of solute

Let mass of water be X

=( 4+x)g

Concentration of a solution (m/m%)= mass of solute/mass of solution× 100

= 4/(4+x)×100= 2

400= 2(4+x)

400= 8+2x

2X= 400-8

2X= 392

X= 196g